# A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images

### v1

@@ VI Conclusion
In scene classification and semantic segmentation, the hierarchical relationship among the target classes is often under-exploited, resulting in the exclusion of valuable semantic information from the decision-making process. To address this limitation, this paper proposes the SAHC approach that leverages the hierarchical structure inherent in class types to improve the robustness and the accuracy of the classification. We introduced hierarchical classification branches, each designed to handle a specific hierarchical classification task, ranging from coarse-grained to fine-grained categories. Trainable hierarchy log-joint matrices are then defined to project the predictions across different hierarchical levels, following a self-supervised data-driven approach which may not align with the rigid user-defined mappings between different hierarchy levels. This allows the implementation of a consensus mechanism, where predictions from multiple hierarchical levels are fused to perform the final decision. By progressively guiding the classification process from broader to specific categories, we can optimize the consensus, ensuring alignment among predictions at different hierarchical levels in a self-supervised fashion. The results show that data-driven hierarchies improve the performance at the finest level compared to directly using the user-defined hierarchies, contrarily to the performance drop usually observed in the literature when adding additional classification layers for different hierarchical levels after the backbone. The proposed methodology is both lightweight and scalable, making it easily adaptable to different architectures and datasets by simply reorganizing their structure in a hierarchical manner.
We evaluated the performance of the proposed approach on three different RS datasets. Indeed, the proposed approach is particularly relevant in VHR image and RS MS SITS classification, where the semantic structure of the data and available datasets support an explicit hierarchical representation of the task. The experimental results reveal that the proposed methodology is able to exploit both complex and simple hierarchies, demonstrating the best performances in terms of OA% and mF1% scores across experiments at the fine-grained hierarchical level. As expected, the improvements are particularly significant when complex hierarchies can be exploited within the RS data. The proposed method is general, showing top performance scores across different tasks and architectures.

### v2

@@ VI Conclusion
In scene classification and semantic segmentation RS tasks, the hierarchical relationship among the target classes is often under-exploited, resulting in the exclusion of valuable semantic information from the decision-making process. To address this limitation, this paper proposes the SAHC approach that leverages the hierarchical structure inherent in class types to improve the robustness and the accuracy of the classification. We introduced hierarchical classification branches, each designed to handle a specific hierarchical classification task, ranging from coarse-grained to fine-grained categories. Trainable hierarchy log-joint matrices are then defined to project the predictions across different hierarchical levels, following a self-consistent data-driven approach which may not align with the rigid user-defined mappings between different hierarchy levels. This allows the implementation of a consensus mechanism, where predictions from multiple hierarchical levels are fused to perform the final decision. By progressively guiding the classification process from broader to specific categories, we can optimize the consensus, ensuring alignment among predictions at different hierarchical levels in a self-consistent fashion. The proposed methodology is both lightweight and scalable, making it easily adaptable to different architectures and datasets by simply reorganizing their structure in a hierarchical manner.
Extensive experimental analysis demonstrated the effectiveness of the proposed approach on two different RS datasets, representing a VHR and a MS classification tasks. Indeed, the proposed approach is particularly relevant in VHR image and RS MS SITS classification, where the semantic structure of the data and available datasets support an explicit hierarchical representation of the task. The experimental results reveal that the proposed methodology is able to exploit both complex and simpler hierarchies, demonstrating the best performances in terms of OA% and mF1% scores across experiments across all hierarchical levels. As expected, the improvements are particularly significant when complex hierarchies can be exploited within the RS data. The proposed method is general, showing top performance scores across different tasks. Furthermore, the self-consistent consensus loss effectively constrained the hierarchical levels to converge toward a unified prediction, reducing hierarchical violations in the final classification maps.